OpenStudy (anonymous):

A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture? If x represents the pounds of $1.50 chocolates used, then which of the following represents the pounds of 90¢ chocolates used? a.x b.30 c.30 - x

Directrix (directrix):

The mixture of the two candies weighs 30 pounds. Let x represent the $1.50 chocolates used. Let y represent the $.90 chocolates used. Then, x + y = 30 -x -x ----------------- y = ? @Jie What y equals is the answer to your question.

OpenStudy (anonymous):

Thank you so much! :)

Directrix (directrix):

Glad to help!